May 4-6

– Three stores provided ~39.75 hours of luminosity
– Tevatron personnel conducted hollow beam studies
– Linac personnel repaired KRF7’s LCW leak
– Operators tuned to correct MI-8 collimator beam positions
– The Thursday day shift was full of accesses for repairs
– MIRF7 tripped off and operators bypassed the station
– Cryo system personnel replaced a wet engine flywheel at F2

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts